OpenStudy (vespersaviour):

A 12-foot ladder is leaning against a wall. The bottom of the ladder is 5 feet away from the bottom of the wall. Approximately how high up the wall does the top of the ladder reach? A. 2.4 feet B. 7.0 feet C. 13.0 feet D. 10.9 feet

OpenStudy (vespersaviour):

I think its C. I dont know if this is correct but 12^2 = 144 5^2 = 25 144 + 25= 169 \[\sqrt{169}=13\] Which is C.

OpenStudy (xxanntdmxx):

a^2 + 5^2 = 12^2 a^2 + 25 = 144 a^2 = 144 - 25 a^2 = 119 a = √119 a ≈ 10.9 feet

OpenStudy (xxanntdmxx):

so im pretty sure its D. :)

OpenStudy (vespersaviour):

Thanks